# An interpretation of scalars in SO(32)

### v1

@@ 1 Introduction
While highly relevant in string theory and supergravity, $SO(32)$ group is not a good unification group as it doesn’t have complex representations [12]. But it stills get an interesting family group when decomposed. In this letter, we first review the decomposition, interpret is as a group symmetry on scalars that could be supersymmetry partners of the Standard Model fermions, and then we present an interesting reconstruction of such scalars as composites. Besides, the interpretation has an uniqueness that limits the number of generations for the SM group.
@@ 2 The flavour group in $SO(32)$
$\begin{array}[]{llrrr}&&Y_{1}&Y_{2}&Q\\ (1,15,\bar{3})&(3,2)&1&-1&+1/3\\ &(6,1)&1&4&-2/3\\ (1,\bar{15},3)&(3,2)&-1&+1&-1/3\\ &(6,1)&-1&-4&+2/3\\ (1,24,1)&(1,1)&0&0&0\\ &(1,3)&0&0&0\\ &(8,1)&0&0&0\\ &(3,2)&0&5&-1\\ &(\bar{3},3)&0&-5&+1\\ \end{array}$ (5)
@@ 3.1 Turtles and elephants
We consider scalars as composites either of pairs of quarks, as a colour triplet, or of pairs quark anti-quark, as a singlet. Furthermore, we divide the quarks in two classes: turtles and elephants, and add a rule: only turtles can combine into composites.
@@ 3.2 Colourless and coloured flavour groups
4 On $SU(15)$
@@ 5 On $SU(8)$
and while in the first approach $SU(5)$ is flavour-colour, in the second it is just two horizontal symmetries and the colour triplet is explicit. So we prefer this later way because so all the irreducible representations of $SU(8)$ have an interesting interpretable descent. The decomposition of the $28$ has a quark content that looks very much as our division in five turtles and one elephant,
@@
but it is different to the $SO(32)$ case. To ilustrate a particular assignment, if we think of the fundamental as ”half-charged preons” of charges ±$\pm 1/2$, $1/6$, then:
@@ 7 Discussion
The postulate It is turtles all the way down [^1] I first heard this idiom in a talk from Alvaro de Rujula in 1986 applied solely to squarks already fixes the number of generations to be greater or equal than three. Adding a reasonable condition on the building of neutral sleptons, it fixes uniquely $N=3$ and then also the separation between five light quarks and one heavy one that does not participate in the composites. Of course this uniqueness is not seen when going directly from the $SO(32)$ group down to flavour times colour, but even in this case there is a separation between five ”turtles” making the fundamental of $SO(32)$ and a non-participant ”elephant”.
While eventually all the extant multiplets of the decomposition should be explained, the $(1,3)$ squarks, of charge ±$\pm 4/3$, are specially puzzling. They can not be organised as three generations of partners of four-component Dirac quarks. Still, they have a role in the flavour multiplet, and they could exhibit their singularity if chirality is introduced back in the game.
The symmetry between quarks and diquarks or its hadronic equivalent is known to be one of the historical origins of supersymmetry [15, 10] and it is used in hadronic phenomenology. But a concrete hadronic construction of our scalars as real diquarks produces the ones of odd parity, that are excluded of phenomenological discussions as they do not survive the ’single mode approximation’ [14]. Thus the composite ”squarks” and ”sleptons” bound here should be not the ones found at QCD scale, but it is intriguing that they are similar in number and mass.
@@
One must recognise that the motivation to use $SO(32)$ is not only to produce one hypercharge and the adequate multiplets in the decomposition, but also because of its role in string theory. The postulates of composition need a pairing that looks similar to labels in terminated open strings. The composition process from the point of view of a terminated ”QCD string” bears some similarity to the techniques of [4] using ”planar orientifolds”.
@@ 8 Conclusions
Our postulate is, certainly, exotic: it suggests that while SSM fermions could be -or not- elementary, the SSM scalars are composites, with their preons being a subset of the fermions. Far fetched as this postulate looks, it reproduces the $SO(32)$ decomposition and fixes the number of possible generations.
@@ 9 Acknowledgements
This work was supported in part by project No. PID2022-136374NB-C22, funded by the Spanish MCIN/AEI and by the European Union. I also thank M Porter for discussions.

### v2

@@ 1 Introduction
While highly relevant in string theory and supergravity, the $SO(32)$ group is not a good unification group as it doesn’t have complex representations [18]. But it still gets an interesting family group when decomposed. In this letter, we first review the decomposition, interpret it as a group symmetry on scalars that could be supersymmetry partners of the Standard Model fermions, and then we present an interesting reconstruction of such scalars as composites. Besides, the interpretation has an uniqueness that limits the number of generations for the SM group.
@@ 2 The flavour group in $SO(32)$
$\begin{array}[]{lcrrr}&(f_{3},f_{2})&Y_{1}&Y_{2}&Q\\ \hline\cr(1,15,\bar{3})&(3,2)&1&-1&+1/3\\ &(6,1)&1&4&-2/3\\ (1,\bar{15},3)&(3,2)&-1&+1&-1/3\\ &(6,1)&-1&-4&+2/3\\ (1,24,1)&(1,1)&0&0&0\\ &(1,3)&0&0&0\\ &(8,1)&0&0&0\\ &(3,2)&0&5&-1\\ &(\bar{3},2)&0&-5&+1\\ \end{array}$ (5)
@@
Figure 1: Illustration of the concept Turtles all way down, with an spectator but massive giant (Credit of the drawing: De Rújula)
@@ 3.1 Turtles and elephants
We consider scalars as composites either of pairs of quarks, as a colour triplet, or of pairs quark anti-quark, as a singlet. Furthermore, we divide the quarks in two classes: turtles and elephants [^1] Or giants, see figure 1, and add a rule: only turtles can combine into composites.
@@ 3.2 Colourless and coloured flavour groups
4 A case for one generation
The above argument assumed the SM and particularly that the turtles were allowed to bind only if they had colour. But the final particle content is very reminiscent of the Georgi-Glashow model.
Consider the usual formulation of the model
$5=$ $(1,2)_{-3}+(3,1)_{2}$ (12)
$10=$ $(1,1)_{-6}+(\bar{3},1)_{4}+(3,2)_{-1}$ (13)
and use $Q=T_{3}-Y/6$ to study the electric charge of the colour singlets and colour triplets in each representation of $SU(5)$
repr singlets triplets antitriplets other
1 0
5 0, +1 -1/3
24 0,+1,0,-1 -4/3, -1/3 +4/3, +1/3 $(8)_{q=0}$
10 +1 +2/3, -1/3 -2/3
$\bar{10}$1 +2/3 -2/3, +1/3
15 +2,+1,0 +2/3, -1/3 $(6)_{-2/3}$
$\bar{15}$2,-1,0 -2/3, +1/3 $(\bar{6})_{+2/3}$
It can be argued that the union of $2410+\bar{10}+$ 24 is the one generation version of our previous construction, albeit with only a left neutrino. See how the $X$ boson of the GUT model is here just the q=+4/3 scalar, and how we have the condition of two states for each fermion. In this case our turtles, binding, are all the members of $5$ and the elephants, not-binding, are the members of $10$. They bypass the previous uniqueness proof because we are allowing leptons to join the game.
It could be interesting to consider variants of this game, such as flipped $SU(5)$ [11] with some rules for the $U(1)_{X}$ charges of the $24$ and $15$, or even an anomalous ”deflipped $SU(5)$” with the up quark in the fundamental but the standard hypercharge assignments, so that the $+4/3$ triplet would appear in the decuplet. The authors in [8] use a different charge assignment to produce the +2/3 charge in the $24$, so that it is composed of gluons, a photon, and a whole set of $SU(3)\times U(1)$ charges, and then a broken susy $SU(5)$ produces the standard model particles as Goldstone fermions.
The $SU(5)$ model was frequently used as foundational model in the peak of composite theories in the early eighties. In most cases the components in the $5$ were new particles, but some models considered to keep right fermion as elementary and only left as composite, and even there was some proposal [13] were the $5$ had a shared mix of preons and known fermions222 Literature reviews also mention ideas where only the leptons are composite but I have not found a concrete reference.
5 Discussion on related groups
5.1 On $SU(15)$
@@ 5.2 On $SU(8)$
and while in the first approach $SU(5)$ is flavour-colour, in the second it is just two horizontal symmetries and the colour triplet is explicit. So we prefer this later way because so all the irreducible representations of $SU(8)$ have an interesting interpretable descent. The decomposition of the $28$ has a quark content, triplets, that looks very much as our division in five turtles and one elephant,
@@
but it is different to the $SO(32)$ case. To illustrate a particular assignment, if we think of the fundamental as ”half-charged preons” of charges ±$\pm 1/2$, $1/6$, then in the $28$:
@@ 6 Discussion
The postulate It is turtles all the way down [^3] I first heard this idiom in a talk from Alvaro de Rujula in 1986 applied solely to squarks already fixes the number of generations to be greater or equal than three. Adding a reasonable condition on the building of neutral sleptons, it fixes uniquely $N=3$ and then also the separation between five light quarks and one heavy one that does not participate in the composites. Of course this uniqueness is not seen when going directly from the $SO(32)$ group down to flavour times colour, but even in this case there is a separation between five ”turtles” in the fundamental of $SO(32)$ and a non-participant ”elephant”.
While eventually all the extant multiplets of the decomposition should be explained, the $(1,3)$ squarks, of charge ±$\pm 4/3$, are specially puzzling. They can not be organised as three generations of partners of four-component Dirac quarks. Still, they have a role in the flavour multiplet, and they could exhibit their singularity if chirality is introduced back in the game. We have not considered other flipped descents that could result in different extra $squarks$
The symmetry between quarks and diquarks or its hadronic equivalent is known to be one of the historical origins of supersymmetry [21, 16] and it is used in hadronic phenomenology. But a concrete hadronic construction of our scalars as real diquarks produces the ones of odd parity, that are excluded of phenomenological discussions as they do not survive the ’single mode approximation’ [20]. Thus the composite ”squarks” and ”sleptons” bound here should be not the ones, diquarks and mesons, found at QCD scale, but it is intriguing that they are similar in number and mass.
Approximate supersymmetry between quarks and diquarks, to be used for dynamical supersymmetry between barions and mesons, is a well known technique, see the review [4]. One of the authors of this review extended this technique to preons [12], so that dipreons are susy partners of preons, but the model does not produces a dynamical susy with new scalars as partners of leptons and quarks.
@@
While the composites point to HC, TC or ETC models, one must recognise that the motivation for $SO(32)$ is not only to produce one hypercharge and the adequate multiplets in the decomposition, but also because of its role in string theory. The postulates of composition need a pairing that looks similar to labels in terminated open strings. The composition process from the point of view of a terminated ”QCD string” bears some similarity to the techniques of [5] using ”planar orientifolds”.
@@ 7 Conclusions
Our postulate is, certainly, exotic: it suggests that while SSM fermions could be -or not- elementary, the SSM scalars are composites, with their preons being a subset of the observed SM fermions. Far fetched as this postulate looks, it reproduces the $SO(32)$ decomposition and fixes the number of possible generations.
@@ 8 Acknowledgements
This work was supported in part by project No. PID2022-136374NB-C22, funded by the Spanish MCIN/AEI and by the European Union. Figure (1) adapted by Álvaro de Rújula from his personal archive. I also thank M Porter for discussions.